The measure of one angle is eleven more than four times a number. Another angle is twice the first angle's measure. The sum of the measures of the angles is 195 degrees. What is the measure of each angle/
Answer by lwsshak3(11628)

The measure of one angle is eleven more than four times a number. Another angle is twice the first angle's measure. The sum of the measures of the angles is 195 degrees. What is the measure of each angle.
***
let x=number
4x+11=measure of 1st angle
2(4x+11)=measure of 2nd angle
4x+11+2(4x+11)=195
4x+11+8x+22=195
12x=162
x=13.5
4x+11=65
2(4x+11=130
measure of 1st angle=65˚
measure of 2nd angle=130˚